Status Report

NASA Hubble Space Telescope Daily Report #4717

By SpaceRef Editor
October 18, 2008
Filed under , ,
NASA Hubble Space Telescope Daily Report #4717


Continuing to collect World Class Science

PERIOD COVERED: 5am October 15 – 5am October 16, 2008 (DOY 289/0900z-290/0900z)


FGS 11212

Filling the Period Gap for Massive Binaries

The current census of binaries among the massive O-type stars is seriously incomplete for systems in the period range from years to millennia because the radial velocity variations are too small and the angular separations too close for easy detection. Here we propose to discover binaries in this observational gap through a Faint Guidance Sensor SNAP survey of relatively bright targets listed in the Galactic O Star Catalog. Our primary goal is to determine the binary frequency among those in the cluster/association, field, and runaway groups. The results will help us assess the role of binaries in massive star formation and in the processes that lead to the ejection of massive stars from their natal clusters. The program will also lead to the identification of new, close binaries that will be targets of long term spectroscopic and high angular resolution observations to determine their masses and distances. The results will also be important for the interpretation of the spectra of suspected and newly identified binary and multiple systems.


Significant Spacecraft Anomalies: (The following are preliminary reports of potential non-nominal performance that will be investigated.)

HSTARS: (None)


18311-2 – Power cycle CU/SDF A @ 289/1102z
18312-0 – Avoid unnecessary FGS FW moves due to default FGE refresh @ 289/1825z
18310-0 – PRT Slew for Post-B-Side Recovery @ 289/1849z
18319-0 – Gyro Bias Update – B-side Switch Step 123 @ 289/1922z
18320-4 – Gyro Bias Estimation OBADs and H&S Intercept @ 289/2118z
18323-0 – Restore NSSC-1/HST486 HV Protection Capability @ 290/0041z
18321-0 – Reinstall Battery VTFE Curves following B-Side Switch @ 290/0058z
18326-0 – Safe and Recover WF/PC-II @ 290/0350z
18325-0 – Omit WFPC2 Dump (Steps 163 & 164) @ 290/0500z
18313-1 – DMU B-Side Switch @ 290/0514z

                       SCHEDULED      SUCCESSFUL

FGS GSacq               05                 05
FGS REacq               00                 00
OBAD with Maneuver      12                 12


1737-0 – LIM OFF for DDIU2AVM @ 289/1246z
1738-0 – PDU2 Current Limit Modification @ 289/1732z
1739-0 – Update CCC V/T Level Yellow and Red Ground Limits @ 289/1740z
1741-0 – Manage ACS WFC Detector Housing Temperature Limit @ 290/0035z
1642-4 – Consolidated PCS Limits @ 290/0149z


Flash Report:

CU/SDF A power cycle confirmed that SDF A is not operable.

Gyro 4 powered up nominally at 289/11:13.

HST was commanded into Hardware Sun-Point safe mode at 8:20 a.m. EDT. The required transitions to the B-sides of some of the Data Management System’s modules proceeded smoothly, and telemetry was reacquired ahead of schedule. The HST486 (spacecraft computer) was reloaded and its performance verified. At 1:10 p.m., HST was commanded out of safe mode. It is again being controlled by the HST486.

All DMS B-Side hardware continues to perform nominally. NICMOS and ACS commanding proceeded nominally and the EEPROM dumps were verified by Payload FSW.

All SIs were safed at 05:10 UTC in preparation for activation of the Science SMS at 290/16:00 UTC. DMS/SI C&DH functions have successfully transitioned to the B-Side and HST is ready for science!

SpaceRef staff editor.